# PocketGS: On-Device Training of 3D Gaussian Splatting for High Perceptual Modeling

### v2

@@ III-C2 Index-Mapped Gradient Scattering
Maintaining optimizer consistency during depth-sorted rendering typically requires physical memory reordering, which incurs heavy data movement. We resolve this by decoupling the sorted view from the canonical layout. A GPU-based sort generates an index list $\pi$, allowing the renderer to access parameters $\theta_{\pi(i)}$ while they remain stationary in memory. During backpropagation, gradients are scattered back to the canonical buffers:
@@
ensuring that Adam optimizer states remain aligned with their respective parameters without CPU intervention or redundant memory copies.
III-C3 On-GPU Parameter Optimization
To eliminate the synchronization latency of CPU-driven updates, the entire optimization loop is executed on-device. We maintain Adam states (moments $m,v$) in dedicated GPU buffers and perform updates via a unified compute kernel. To ensure convergence under limited FP16 precision, we employ specialized parameterization: Opacity uses logit-space updates with a micro-step threshold; Scale is optimized in log-space with explicit clamping; and Rotation is updated via tangent-space projection and sign-unification. This fully GPU-resident schedule minimizes the host-device communication footprint while maximizing hardware utilization.
@@ V Conclusion
We introduce PocketGS, a fully on-device 3DGS training paradigm that co-designs geometry priors ($\mathcal{G}$), prior-conditioned initialization ($\mathcal{I}$), and hardware-aligned splatting optimization ($\mathcal{T}$) to make high-fidelity 3D reconstruction practical under mobile compute and memory constraints. PocketGS achieves perceptual quality competitive with or superior to workstation baselines while meeting smartphone budgets (e.g., $\sim$4-minute end-to-end training and <3<3 GB peak memory on iPhone 15), enabling practical professional-grade mobile 3D content creation for mixed reality and digital twins.

### v3

@@ III-C2 Index-Mapped Gradient Scattering
Maintaining optimizer consistency during depth-sorted rendering typically requires physically reordering parameter buffers, which incurs heavy data movement and can misalign optimizer states. We resolve this by decoupling the sorted view used for rendering/backprop from the canonical layout used for parameter storage and optimizer moments. A GPU-based sort produces a permutation $\pi$ that maps each sorted index $i$ to its canonical index $\pi(i)$. The renderer and backward kernels operate on the sorted view, while gradients are scattered back to canonical buffers:
@@
ensuring that Adam optimizer states (stored in canonical order) remain aligned with their respective parameters, without CPU intervention or redundant memory copies.
@@ V Conclusion
In this paper, we present PocketGS, a fully on-device training framework for 3D Gaussian Splatting that targets the practical constraints of mobile devices. PocketGS is built upon three co-designed operators: (G) geometry-prior construction to provide compact and reliable geometric guidance, (I) prior-conditioned Gaussian parameterization to improve initialization and optimization under iteration-limited budgets, and (T) hardware-aligned splatting optimization to ensure efficient and stable differentiable splatting on mobile GPUs. Extensive experiments on public benchmarks and our MobileScan dataset show that PocketGS achieves strong perceptual quality while meeting smartphone budgets, e.g., approximately 4-minute end-to-end training with <3<3 GB peak memory on iPhone 15. These results indicate that high-fidelity 3D reconstruction and mobile 3D content creation can be made practical on commodity devices.